Solve the following recurrence relation




a) T(n)=T(n-1) +1,for n>2 and T(1)=1

1
Expert's answer
2022-05-03T10:00:19-0400

T(1)=1

T(2)=1+1=2

T(3)=2+1=3

T(4)=3+1=4

T(n)=n